# Zhuoming Chen Ranajoy Sadhukhan Zihao Ye Yang Zhou Jianyu Zhang Niklas Nolte

### v2

@@ 4.1 Self-normalized importance sampling for attention estimation
Fortunately, importance sampling [38, 48, 45] allows us to estimate unknown distribution $w$ by sampling from a proposed distribution $u$. In our problem setting, the normalization factor of $w$, i.e. $Z=\sum_{i=1}^{n}\exp{\frac{qk_{i}^{T}}{\sqrt{d}}}$ is also unknown because computing it requires evaluating all $qk_{i}^{T}$. However, we do have access to unnormalized weights $\widetilde{w_{i}}=e^{\frac{qk_{i}^{T}}{\sqrt{d}}}$ for sampled indices $i$. Hence, by employing a variant of importance sampling, self-normalized importance sampling [48], we sample indices $i_{1},i_{2},...,i_{\mathcal{B}}$ from a proposed distribution $u$ and the resulting estimator is
@@
which has a very nice property for accurately estimating attention output that $\mathbb{P}[\lim_{k\rightarrow\infty}X^{\mathrm{IS}}=o]=1$. Its variance [^1] We assume head dimension $d=1$ here for simplicity. Higher dimensions have similar formulations and analyses by replacing variance with the trace of covariance. is related to the distribution $u$, and can be approximated by
@@
To minimize the variance, $u$ should satisfy $u\propto\widetilde{w_{i}}|v_{i}-o|$ [30]. The variance will be high if $u_{i}$ and $\widetilde{w_{i}}|v_{i}-o|$ assign a high probability mass to different regions of the sample space or have different modes. Therefore, the challenge is computing a distribution $u$ aligned with $\widetilde{w_{i}}|v_{i}-o|$ without accessing too many $\widetilde{w_{i}}$. Besides, Equation 6 requires that sampling probability $u$ can be computed and $u_{i}>0$, which is not satisfied by many deterministic approximations like $\mathrm{TopK}$.
@@ 4.2 Variance reduction with LSH
we further transfer the inner product $qk_{i}^{T}$ to cosine similarity between $\bar{q}$ and $\bar{k_{i}}$ (which is a common practice in Maximum Inner Product Search [54]).
Inspired by prior work [56, 14], we leverage Locality sensitive hashing-based sampling for this estimation problem. Specifically, leveraging a hash function $h$ in the LSH family that preserves cosine similarity such as SimHash [53], we can sample from probability distribution $u_{i}=\mathbb{P}[{h}(q)=h(k_{i})]$ which is monotonic to $\cos{\frac{qk_{i}^{T}}{|q|\cdot|k_{i}|}}$.
4.3 Algorithm implementation
@@
Hash function selection. MagicPIG leverages SimHash [53], that draws with $K\times L$ random vectors. For each of the $L$ hash tables, the ${q}$ and ${k_{i}}$s vectors are projected on $K$ directions, and only the sign of the projection is kept, which yields a $K$-bit hash value. Key ${k_{i}}$ is sampled only if there exist at least two hash tables where ${k_{i}}$ shares the hash value with ${q}$. The corresponding probability is
@@
Input: $\bm{K},\bm{V}\in{R}^{n\times d}$, $\bm{q}\in{R}^{1\times d}$, random projectors $\bm{W}\in R^{d\times(K\times L)}$, hash tables $\bm{HT}$, static KV cache $\bm{K_{T}},\bm{V_{T}}\in{R}^{t\times d}$.
@@
Query hash tables to sample $\bm{S}$ in Equation 9
$\bm{S}$ = $\textbf{Query}(\bm{HT},\bm{q}_{\text{code}}),\bm{K_{S}}=\bm{K[S]},\bm{V_{S}}=\bm{V[S]}$
Compute inner product for $\bm{q}$ and sampled $\bm{K}$
$\bm{w_{S}}=\bm{q}\bm{K^{T}_{S}}$, $\bm{w_{T}}=\bm{q}\bm{K^{T}_{T}}$
@@
$\bm{p}=1-{\bm{w}}/{(||\bm{q}||\cdot||\bm{K_{S}}||})$
@@
$\bm{\bar{o}}=\textbf{Softmax}(\frac{[\bm{w_{S}},\bm{w_{T}}]}{\sqrt{d}}-\log([\bm{u},\bm{1_{t}}]))[\bm{V_{S}},\bm{V_{T}}]$
Return $\bm{\bar{o}}$
@@
Data pre-processing. Before building hash tables, MagicPIG centers the $k_{i}$ vectors. As shown in Figure 2c, keys are almost always concentrated on one side of the queries, except the initial token. In this case, random projections cannot effectively distinguish keys, resulting in uniform sampled probabilities. $\mathrm{Softmax}$ is translation invariant. Centering ($\bar{k_{i}}=k_{i}-\frac{1}{n}\sum_{i=1}^{n}k_{i}$) distributed the keys better and remains computationally equivalent.
@@ 4.4 System co-design
The memory size of KV cache remains a bottleneck for long-context LLM decoding, especially when GPU VRAM is limited. DRAM on the CPU side offers sufficient memory capacity with $100-200$GB/s bandwidth, which is usually $10-20\%$ of GPU VRAM bandwidth (see Figure 7a). Ideally, this gap can be mitigated by 510×5-10$\times$5 - 10 × sparsity. To make CPU DRAM an aggregated memory for GPU, the workload must be partitioned. In our experiments, $K=9$ or 10, and $L$ is a few hundred.
Our system design extends prior work [27, 5] by splitting LLM decoding into three parts. (1) Parameter computations, i.e., all linear projectors including MLP and $W_{Q},W_{K},W_{V},andW_{O}$ in the self-attention module run on GPU. (2) Attention computation, which involves $o=\mathrm{Softmax}(\frac{qK^{T}}{\sqrt{d}})V$, runs on CPU. (3) Random projections. At generation time, for each $q$, $K\times L$ random projections are conducted to obtain the hash codes. Since all heads can share the same random projectors, the memory overhead is limited (25 MB in our implementation), so this step is compute-bound. Therefore, the projection is placed on GPU. (4) Retrieval. The hash codes of $q$, need to be looked up in $L$ hash tables, which is negligible computationally. However, the pre-built hash tables for $k_{i}$s can occupy considerable memory, making it a better fit for the CPU. With the above partition, we are able to support hash tables with $K$ and $L$ beyond the scale of prior work [37, 16, 65] without worrying about computation for hash codes as well as the storage of hash tables.
On-device cache. Sink tokens (the first several tokens) and local tokens are more likely to be sampled according to their high similarity to the query. To further reduce CPU workload, MagicPIG stores these tokens on GPU and does not apply LSH sampling to them. We leverage the recursive attention technique [64] to merge the attention output from CPU and GPU.
Our algorithm applies to a single attention head, see Algorithm 1. The details of Encode, Query, as well as the hash table construction, are described in prior work [53, 15].
@@ 5 Evaluation
  • In Figure 8, we demonstrate the system performance of MagicPIG, which achieves up to 3.9×3.9$\times$3.9 × throughput improvement and 110ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct with 96K context.
@@ 5.1 MagicPIG Preserves Accuracy
Setup. Our experiments are based on Llama [3, 25, 60] models. Three types of tasks are included, which are 3 mid-context comprehensive tasks from lm-eval-harness [26] (GSM8K-CoT [19], MMLU-Flan-Cot-Fewshot [29] and COQA [51]), and 6 long context tasks from [8] (QASPER [23], LCC, Repobench-P [43], TriviaQA [35], PRE and TREC [39, 32]) and 13 synthetic tasks from RULER [33] (with 50 examples per task).
Baselines. Besides full attention, Quest [58] and its variants are used as baselines. In its default setting, Quest uses a “page size” of 16, i.e. 1/16 of the full attention cost. To compare the methods fairly in the low computation budget regime, we also evaluate Quest with page size 32 and 64 and make sure at least one page is selected in every test example. The initial 4 tokens and local 64 (for LongBench [8] and RULER [33]) or 24 (for lm-eval-harness [26]) tokens as well as layer-$\{0,16\}$ are statically preserved. We do not use the theoretical transformations in Equation 8 in our implementations, as we do not find them to contribute to accuracy improvements.
@@
Analysis. From Tables 1, 2 and 3, (1) MagicPIG preserves high accuracy (degradation less than $2\%$) for all kinds of tasks, with a computation cost of $2\%\sim 5\%$. (2) Compared with Quest, which also shows reasonable performance on long context tasks, MagicPIG also demonstrates good performance on tasks with moderate context sizes in lm-eval-harness [26], indicating a more robust performance in general serving. (3) With LSH sampling, which introduces an order of magnitude lower sampling/searching cost ($\text{Cost}_{1}$), MagicPIG can achieve equivalent or better accuracy with only half of the computation cost.
Table 1: Comprehensive tasks on lm-eval-harness [26]. MagicPIG significantly outperforms other methods with lower computation. The config (K, L) is a hyper-parameter of LSH for MagicPIG or page size and ratio of selected pages for Quest [58]. $\text{Cost}_{1}$, $\text{Cost}_{2}$ represents the cost for searching/sampling and sparse attention computation.
@@
Table 2: Long context tasks on LongBench [8]. MagicPIG preserves high accuracy with low computation. Config and cost are defined as in Table 1. Code models are only evaluated by Repobench-P and LCC.
@@
Table 3: Synthesized tasks on RULER [33]. MagicPIG preserves high accuracy with low computation. Config and cost are defined as in Table 1.
@@ 5.2 MagicPIG Shows Impressive Efficiency across Various Hardware Settings
We show MagicPIG can bring up to 3.9×3.9$\times$3.9 × wall clock speed up and reduce GPU memory consumption on different models and hardware settings (A100, L40, RTX4090).
(a) A100 with 34B model
(b) L40 with 13B model
(c) RTX 4090 with 8B model
Figure 8: We evaluate MagicPIG on three serving scenarios. Left: A100 serves 34B model with 16K context. MagicPIG achieves 1.9×1.9$\times$1.9 × throughput improvement. Mid: L40 serves 13B model with 16K context. MagicPIG achieves 3.9×3.9$\times$3.9 × throughput improvement. Right: Simulated RTX 4090 serves 8B model with 128K context. MagicPIG achieves a latency of 110ms in a single request serving and can improve the throughput of baseline by up to 2.9×2.9$\times$2.9 ×. The dashed lines denote the highest throughput of baselines. With KV cache offloading, MagicPIG can fit a much larger batch size compared with full attention on GPU, which contributes to the throughput improvement.
Setup. We evaluate our system performance on 3 serving settings. (1) 80GB GPU (A100) and 34B model (CodeLlama-34B) [52] with 16K contexts; (2) 48GB GPU (L40) and 13B model (CodeLlama-13B) [52] with 16K contexts; (3) 24GB GPU [^2] We simulate 24GB GPU by setting memory limit with L40. As the bandwidth of L40 (864GB/s) is less than RTX 4090 (1TB/s), the real speed of our system should be slightly faster than the simulation. (e.g. RTX 4090) and 8B model (Llama-3.1-8B) [25] with 96K contexts.
Baselines. Our baselines for (1) and (2) are full attention on GPU, and for (3) is full attention on CPU with theoretical estimated bandwidth. Our system’s GPU part is implemented in native Pytorch [49] and the CPU part in FBGEMM [36] in bfloat16 precision. Our CPU is Intel(R) Xeon(R) Platinum 8480+ for A100 and Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz for L40. In the last setting, the CPU bandwidth is estimated at 100GB/s, above the empirical bandwidth we measure when running a group query attention of size 4.
Analysis. In Figures 8a, 8c and 8b, we demonstrate (1) MagicPIG significantly improves decoding throughput for all three scenarios (A100: 1.9×1.9$\times$1.9 ×, L40: 3.9×3.9$\times$3.9 ×, RTX 4090: 2.9×2.9$\times$2.9 ×) and can achieve a latency of 110ms for single request generation with 96K context for RTX 4090. (2) With KV cache offloading, MagicPIG can fit much larger batches than GPU full attention baselines (1518×15$\sim$ 18$\times$15 ∼ 18 ×).

### v4

@@ 4.1 Self-normalized importance sampling for attention estimation
Fortunately, importance sampling (Kloek and Van Dijk, 1978; Owen, 2013; Lohr, 2021) allows us to estimate unknown distribution $w$ by sampling from a proposed distribution $u$. In our problem setting, the normalization factor of $w$, i.e. $Z=\sum_{i=1}^{n}\exp{\frac{qk_{i}^{T}}{\sqrt{d}}}$ is also unknown because computing it requires evaluating all $qk_{i}^{T}$. However, we do have access to unnormalized weights $\widetilde{w_{i}}=e^{\frac{qk_{i}^{T}}{\sqrt{d}}}$ for sampled indices $i$. Hence, by employing a variant of importance sampling, self-normalized importance sampling (Owen, 2013), we sample indices $i_{1},i_{2},...,i_{\mathcal{B}}$ from a proposed distribution $u$ and the resulting estimator is
@@
which has a very nice property for accurately estimating attention output that $\mathbb{P}[\lim_{\mathcal{B}\rightarrow\infty}X^{\mathrm{IS}}=o]=1$.
Its variance [^1] We assume head dimension $d=1$ here for simplicity. Higher dimensions have similar formulations and analyses by replacing variance with the trace of covariance. is related to the distribution $u$, and can be approximated by
@@
To minimize the variance, $u$ should satisfy $u\propto\widetilde{w_{i}}|v_{i}-o|$ (Hesterberg, 2003). The variance will be high if $u_{i}$ and $\widetilde{w_{i}}|v_{i}-o|$ assign a high probability mass to different regions of the sample space or have different modes. Therefore, the challenge is computing a distribution $u$ aligned with $\widetilde{w_{i}}|v_{i}-o|$ without accessing too many $\widetilde{w_{i}}$. Besides, Equation 6 requires that sampling probability $u$ can be computed and $u_{i}>0$, which is not satisfied by many deterministic approximations like $\mathrm{TopK}$.
@@ 4.2 Variance reduction with LSH
we further transfer the inner product $qk_{i}^{T}$ to cosine similarity between $\bar{q}$ and $\bar{k_{i}}$ (which is a common practice in Maximum Inner Product Search (Shrivastava and Li, 2014)).
Inspired by prior work (Spring and Shrivastava, 2017; Chen et al., 2020a), we leverage Locality sensitive hashing-based sampling for this estimation problem. Specifically, leveraging a hash function $h$ in the LSH family that preserves cosine similarity such as SimHash (Sadowski, 2007), we can sample from probability distribution $u_{i}=\mathbb{P}[{h}(q)=h(k_{i})]$ which is monotonic to $\cos{\frac{qk_{i}^{T}}{|q|\cdot|k_{i}|}}$.
4.3 Algorithm Design
@@
Hash function selection. MagicPIG leverages SimHash (Sadowski, 2007), that draws with $K\times L$ random vectors. For each of the $L$ hash tables, the ${q}$ and ${k_{i}}$s vectors are projected on $K$ directions, and only the sign of the projection is kept, which yields a $K$-bit hash value. Key ${k_{i}}$ is sampled only if there exist at least two [^2] Empirical results show that requiring hits in two hash tables greatly improves accuracy over standard SimHash. hash tables where ${k_{i}}$ shares the hash value with ${q}$. The corresponding probability is
@@
Input: $\boldsymbol{K},\boldsymbol{V}\in{R}^{n\times d}$, $\boldsymbol{q}\in{R}^{1\times d}$, random projectors $\boldsymbol{W}\in R^{d\times(K\times L)}$, hash tables $\boldsymbol{HT}$, static KV cache $\boldsymbol{K_{T}},\boldsymbol{V_{T}}\in{R}^{t\times d}$.
@@
Query hash tables to sample $\boldsymbol{S}$ in Equation 9
$\boldsymbol{S}$ = $\textbf{Query}(\boldsymbol{HT},\boldsymbol{q}_{\text{code}}),\boldsymbol{K_{S}}=\boldsymbol{K[S]},\boldsymbol{V_{S}}=\boldsymbol{V[S]}$
Compute inner product for $\boldsymbol{q}$ and sampled $\boldsymbol{K}$
$\boldsymbol{w_{S}}=\boldsymbol{q}\boldsymbol{K^{T}_{S}}$, $\boldsymbol{w_{T}}=\boldsymbol{q}\boldsymbol{K^{T}_{T}}$
@@
$\boldsymbol{p}=1-\frac{1}{\pi}\arccos({{\boldsymbol{w}}/{(||\boldsymbol{q}||\cdot||\boldsymbol{K_{S}}||})})$
@@
$\boldsymbol{\bar{o}}=\textbf{Softmax}(\frac{[\boldsymbol{w_{S}},\boldsymbol{w_{T}}]}{\sqrt{d}}-\log([\boldsymbol{u},\boldsymbol{1_{t}}]))[\boldsymbol{V_{S}},\boldsymbol{V_{T}}]$
Return $\boldsymbol{\bar{o}}$
@@
Data pre-processing. Before building hash tables, MagicPIG centers the $k_{i}$ vectors. As shown in Figure 2(c), keys are almost always concentrated on one side of the queries, except the initial token. In this case, random projections cannot effectively distinguish keys, resulting in uniform sampled probabilities. $\mathrm{Softmax}$ is translation invariant. Centering ($\bar{k_{i}}=k_{i}-\frac{1}{n}\sum_{i=1}^{n}k_{i}$) distributed the keys better and remains computationally equivalent.
@@ 4.4 System co-design
The memory size of KV cache remains a bottleneck for long-context LLM decoding, especially when GPU VRAM is limited. DRAM on the CPU side offers sufficient memory capacity with $100-200$GB/s bandwidth, which is usually $10-20\%$ of GPU VRAM bandwidth (see Figure 7(a)). Ideally, this gap can be mitigated by 510×5-10$\times$5 - 10 × sparsity. To make CPU DRAM an aggregated memory for GPU, the workload must be partitioned. In our experiments, $K=9$ or 10, and $L$ is a few hundred.
Our system design extends prior work (He and Zhai, 2024; Aminabadi et al., 2022) by splitting LLM decoding into three parts. (1) Parameter computations, i.e., all linear projectors including MLP and $W_{Q},W_{K},W_{V},andW_{O}$ in the self-attention module run on GPU. (2) Attention computation, which involves $o=\mathrm{Softmax}(\frac{qK^{T}}{\sqrt{d}})V$, runs on CPU. (3) Random projections. At generation time, for each $q$, $K\times L$ random projections are conducted to obtain the hash codes. Since all heads can share the same random projectors, the memory overhead is limited (400 KB in our implementation), so this step is compute-bound. Therefore, the projection is placed on GPU. (4) Retrieval. The hash codes of $q$, need to be looked up in $L$ hash tables, which is negligible computationally. However, the pre-built hash tables for $k_{i}$s can occupy considerable memory, making it a better fit for the CPU. With the above partition, we are able to support hash tables with $K$ and $L$ beyond the scale of prior work (Kitaev et al., 2020; Chen et al., 2021; Zandieh et al., 2023) without worrying about computation for hash codes as well as the storage of hash tables.
On-device cache. Sink tokens (the first several tokens) and local tokens are more likely to be sampled according to their high similarity to the query. To further reduce CPU workload, MagicPIG stores these tokens on GPU and does not apply LSH sampling to them. We leverage the recursive attention technique (Ye et al., 2024) to merge the attention output from CPU and GPU.
Our algorithm applies to a single attention head, see Algorithm 1. The details of Encode, Query, as well as the hash table construction, are described in prior work (Sadowski, 2007; Chen et al., 2020b). In Section 11, we discuss the selection of LSH hyper-parameter (K, L).
@@ 5 Evaluation
  • In Figure 8, we demonstrate the system performance of MagicPIG, which achieves up to 5×5$\times$5 × throughput improvement and 54ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct with 96K context.
@@ 5.1 MagicPIG Preserves Accuracy
Setup. Our experiments are based on Llama (AI@Meta, 2024; Dubey et al., 2024; Touvron et al., 2023) models. Three types of tasks are included, which are 3 mid-context comprehensive tasks from lm-eval-harness (Gao et al., 2021) (GSM8K-CoT (Cobbe et al., 2021), MMLU-Flan-Cot-Fewshot (Hendrycks et al., 2020) and COQA (Reddy et al., 2019)), and 6 long context tasks from (Bai et al., 2023) (QASPER (Dasigi et al., 2021), LCC, Repobench-P (Liu et al., 2023), TriviaQA (Joshi et al., 2017), PRE and TREC (Li and Roth, 2002; Hovy et al., 2001)) and 13 synthetic tasks from RULER (Hsieh et al., 2024) (with 50 examples per task).
Baselines. Besides full attention, Quest (Tang et al., 2024) and its variants are used as baselines. In its default setting, Quest uses a “page size” of 16, i.e. 1/16 of the full attention cost. To compare the methods fairly in the low computation budget regime, we also evaluate Quest with page size 32 and 64 and make sure at least one page is selected in every test example. The initial 4 tokens and local 64 (for LongBench (Bai et al., 2023) and RULER (Hsieh et al., 2024)) or 24 (for lm-eval-harness (Gao et al., 2021)) tokens as well as layer-$\{0,16\}$ are statically preserved. We do not use the theoretical transformations in Equation 8 in our implementations, as we do not find them to contribute to accuracy improvements.
@@
Analysis. From Tables 1, 2 and 3, (1) MagicPIG preserves high accuracy (degradation less than $2\%$) for all kinds of tasks, with a computation cost of $2\%\sim 5\%$. (2) Compared with Quest, which also shows reasonable performance on long context tasks, MagicPIG also demonstrates good performance on tasks with moderate context sizes in lm-eval-harness (Gao et al., 2021), indicating a more robust performance in general serving. (3) With LSH sampling, which introduces an order of magnitude lower sampling/searching cost ($\text{Cost}_{1}$), MagicPIG can achieve equivalent or better accuracy with only half of the computation cost.
Table 1: Comprehensive tasks on lm-eval-harness (Gao et al., 2021). MagicPIG significantly outperforms other methods with lower computation. The config (K, L) is a hyper-parameter of LSH for MagicPIG or page size and ratio of selected pages for Quest (Tang et al., 2024). $\text{Cost}_{1}$, $\text{Cost}_{2}$ represents the cost for searching/sampling and sparse attention computation.
@@
Table 2: Long context tasks on LongBench (Bai et al., 2023). MagicPIG preserves high accuracy with low computation. Config and cost are defined as in Table 1. Code models are only evaluated by Repobench-P and LCC.
@@
Table 3: Synthesized tasks on RULER (Hsieh et al., 2024). MagicPIG preserves high accuracy with low computation. Config and cost are defined as in Table 1.
@@
Llama3-8B-Prolong-512K Full 93.5 90.8 85.1 83.5 88.2 0.00 1.00 1.00
MagicPIG (10,150) 88.0 86.4 81.3 78.8 83.6 0.00 0.02 0.02
MagicPIG (10,170) 89.0 88.7 82.8 80.0 85.1 0.00 0.025 0.025
MagicPIG (9,120) 91.4 88.2 82.4 80.4 85.6 0.00 0.04 0.04
MagicPIG (8,75) 91.4 88.6 83.1 80.5 85.9 0.00 0.05 0.05
Quest (16,0.04) 84.9 83.7 78.7 78.6 81.5 0.06 0.04 0.10
@@ 5.2 MagicPIG Shows Impressive Efficiency across Various Hardware Settings
We show MagicPIG can bring up to 5×5$\times$5 × wall clock speed up and reduce GPU memory consumption on different models and hardware settings (A100, L20, RTX4090).
Figure 8: We evaluate MagicPIG on three serving scenarios. Left: A100 serves 34B model with 16K context. MagicPIG achieves 1.5×1.5$\times$1.5 × throughput improvement. Mid: L20 serves 13B model with 16K context. MagicPIG achieves 5.0×5.0$\times$5.0 × throughput improvement. Right: Simulated RTX 4090 serves 8B model with 96K context. MagicPIG achieves a latency of 54ms in a single request serving and can improve the throughput of baseline by up to $\times3$. The dashed lines denote the highest throughput of baselines. With KV cache offloading, MagicPIG can fit a much larger batch size compared with full attention on GPU, which contributes to the throughput improvement.
Setup. We evaluate our system performance on 3 serving settings. (1) 80GB GPU (A100) and 34B model (CodeLlama-34B) (Rozière et al., 2024) with 16K contexts; (2) 48GB GPU (L20) and 13B model (CodeLlama-13B) (Rozière et al., 2024) with 16K contexts; (3) 24GB GPU [^3] We simulate 24GB GPU by setting memory limit with L20. As the bandwidth of L20 (864GB/s) is less than RTX 4090 (1TB/s), the real speed of our system should be slightly faster than the simulation. (e.g. RTX 4090) and 8B model (Llama-3.1-8B) (Dubey et al., 2024) with 96K contexts.
Baselines. Our baselines for (1) and (2) are full attention on GPU, and for (3) is full attention on CPU with theoretical estimated bandwidth. Our system’s GPU part is implemented in native Pytorch (Paszke et al., 2019) and the CPU part in FBGEMM (Khudia et al., 2021) in bfloat16 precision. Our CPU is Intel Platinum 8480+ for A100 and Intel 8563C for L20. In the last setting, the CPU bandwidth is estimated at 150GB/s, above the empirical bandwidth we measure when running a group query attention of size 4.
Analysis. In Figures 8(a), 8(c) and 8(b), we demonstrate (1) MagicPIG significantly improves decoding throughput for all three scenarios (A100: 1.5×1.5$\times$1.5 ×, L20: 5.0×5.0$\times$5.0 ×, RTX 4090: $\times3$) and can achieve a latency of 54ms for single request generation with 96K context for RTX 4090. (2) With KV cache offloading, MagicPIG can fit much larger batches than GPU full attention baselines (over 12×12$\times$12 ×). The ablation study of decoding throughput with different LSH hyper-parameters is presented in Table 7.